### v2

# On the Element-Wise Representation and Reasoning in Zero-Shot Image Recognition: A Systematic Survey

Abstract
Zero-shot image recognition (ZSIR) aims at empowering models to recognize and reason in unseen domains via learning generalized knowledge from limited data in the seen domain. The gist for ZSIR is to execute element-wise representation and reasoning from the input visual space to the target semantic space, which is a bottom-up modeling paradigm inspired by the process by which humans observe the world, i.e., capturing new concepts by learning and combining the basic components or shared characteristics. In recent years, element-wise learning techniques have seen significant progress in ZSIR as well as widespread application. However, to the best of our knowledge, there remains a lack of a systematic overview of this topic. To enrich the literature and provide a sound basis for its future development, this paper presents a broad review of recent advances in element-wise ZSIR. Concretely, we first attempt to integrate the three basic ZSIR tasks of object recognition, compositional recognition, and foundation model-based open-world recognition into a unified element-wise perspective and provide a detailed taxonomy and analysis of the main research approaches. Then, we collect and summarize some key information and benchmarks, such as detailed technical implementations and common datasets. Finally, we sketch out the wide range of its related applications, discuss vital challenges, and suggest potential future directions.
Index Terms:
Zero-Shot Learning, Pattern Recognition, Representation Learning, Image Processing, Computer Vision.
I Introduction
With the gradual shift from traditional machine learning models to deep neural networks, automated image recognition tasks are gradually achieving accuracy comparable to that of the human brain. The success of deep networks cannot be achieved without two main factors, i.e., massive data along with its annotation and a closed-form environment. The deficiencies are obvious. On the one hand, the cost of image collection is unaffordable; ImageNet [1] contains 14 million images with only 21,814 categories, while most species are much more than that, such as a conservative estimate that there are at least 3 million insect species in the world. On the other hand, image samples of some data are unavailable, such as rare plants, rare medical cases and privacy-protected data. In addition, the requirement for automated models has become more complex now, not only to know what an object is, but also to describe its state [2, 3], e.g., the color of a car, the age of a dog, and so on. The superposition of states and objects means that the cost and difficulty of data collection increase geometrically. As a result, in recent years the interest of researchers has gradually moved from closed-set accuracy to open environment deployment, whose central claim is to endow the model with the ability to discriminate new things without the need to see any visual samples in advance [4, 5, 6]. The way humans think makes the aspiration possible. People can recognize things by extracting and reorganizing their shared characteristics for logical reasoning [7]. For example, we can identify a zebra by combining the morphology of a horse, the stripes of a tiger, and the color of a panda without ever having seen an image of it, and similarly identify a blue car and a red bicycle by combining a red car and a blue bicycle. It implies that human cognition of upper things can be translated into the dismantling and reconstruction of base elements. We have to ask: Can this paradigm of mind be transplanted to the modern automatic recognition system? Naturally, this question motivates the development and application of element-wise learning in open-world recognition tasks.
The objective of element-wise learning is to construct a bottom-up cognitive system by executing element-wise representation and reasoning, i.e., a pyramidal knowledge architecture from micro to macro, attribute to object, component to whole, and local to global. Elements are carriers of properties [8], which describe the meta concepts of a thing, such as materials, colors, patterns, styles, expressions, parts, or functions. The semantic information conveyed by an element varies depending on the specific task and scene. It can be concrete components, e.g., wings and claws, in bird recognition [9] or abstract descriptions, e.g., studying and swimming, in scene classification [10]. Employing element-wise representation to capture discriminative semantic information is already customary for fine-grained classification tasks [11]. And it has been proven effective in many other areas, such as face recognition [12, 13], image retrieval [14] and visual question answering [15]. Element-wise reasoning is the ability to process and apply a complex set of elements. The crux of the extension to the open environment is the combination of different elements to learn new concepts. It is inspired by the hallmark of human intelligence [3], which possesses systematic and structured combinational logic such as predicate logic, boolean logic, and so on [16]. From an early stage, researchers have explored the relationship between combinational elements and objects in visual recognition [17, 18]. Many modern technologies are also designed to follow the idea of combinational reasoning, such as deformable network [19] and graph [7, 20].
Element-wise learning and Zero-Shot Image Recognition (ZSIR) are highly compatible, and the deep integration of the two provides a powerful solution for open-environment deployment. ZSIR employs text as a medium to bridge visual information between known and unknown classes, thus generalizing the model to the unseen domain. Simply put, the shared visual information is compressed into semantic text, and the model unifies the semantic text and image features of the known classes to reason about the unknown classes. As one of the many illuminating solutions, ZSIR has unique advantages. Comparing to other low-shot schemes such as few-shot learning [21] and one-shot learning [22], ZSIR reduces the data requirement to 00 thus completely releasing the data pressure. Comparing to Out-Of-Distribution detection [23], ZSIR assigns precise labels for unknown samples instead of simply categorizing them as unknown. Element-wise learning provides an effective guideline for ZSIR, which profoundly mimics the way humans observe the world, learning the basics through fine-grained language-vision interactions to identify new concepts.
In recent years, ZSIR techniques based on element-wise representation and reasoning have been rapidly developed and widely applied. Unfortunately, the related studies are fragmented and split, lacking a unified and systematic investigation. Although some surveys [4, 24, 6] have been presented, they do not delve in-depth into element-wise techniques. Worse still, they focus only on zero-shot object recognition, which is one of the subtasks of ZSIR, and thus are far from sufficient to show the holistic portrait of ZSIR. To this end, we conduct an extensive survey to thoroughly explore the current situation of the development of element-wise ZSIR techniques. And we attempt to integrate the three mainstream tasks of ZSIR, including object recognition, compositional recognition, and foundation model-based open-world recognition, to provide a more comprehensive view.
Our main contributions are summarized below:
  • We integrate three major zero-shot image recognition tasks, including object recognition, compositional recognition, and foundation model-based open-world recognition, to provide a comprehensive view of element-wise zero-shot learning.
  • We discuss the differences and synergies of several tasks, and analyze the key challenges. We give a detailed introduction to the cutting-edge research techniques and provide a new taxonomy.
  • We collect and present some key information, such as technical details and public datasets.
  • we present the wide range of applications and discuss the potential future directions.
II Description, Comparison, Challenge, and Organization
II-A Description
Figure 1: Three major masks of zero-shot image recognition. Zero-Shot Object Recognition utilizes attributes shared between classes to identify unseen samples. Zero-Shot Compositional Recognition reasons about unseen compositions by learning seen compositions. Foundation model-Based Open-world Recognition exploits the broad fundamentals learned by pre-trained VLMs to implement zero-shot recognition directly in downstream tasks.
Existing surveys only address the topic of zero-shot object recognition, which is far from covering the entire research of ZSIR. In contrast, we integrate three mainstream tasks to provide a more comprehensive view, including object recognition, combination recognition, and foundation model-based open-world recognition. Firstly, we give a detailed introduction to the three tasks.
Zero-Shot Object Recognition (ZSOR). A plain desire is to learn classification in the absence of images to alleviate the dilemma of unavailability of visual data. Humans can recognize 30,000 categories even if some of them have never been seen before [17]. It relies on the full utilization of both verbal and visual media, where we can ”imagine” what an image looks like through its description. Zero-Shot Learning (ZSL) [25] is born to explore the possibility of realizing this capability on automated models. ZSL provides attribute annotations describing the categories, which may be words, sentences, or semantic embeddings. In the training phase, only attribute annotations and images of partial categories are available. The model needs to recognize those categories that have never been seen before in the inference phase. Therefore, the attribute annotations of unseen categories naturally serve as a nexus for conveying the semantics. The setting of ZSL is a bit limited as it only requires testing unseen categories, which is clearly not consistent with real scenarios. Generalized Zero-Shot Learning (GZSL) [4] further extends the setting by requiring both unseen and seen categories to be recognized during the testing period. GZSL is more practical and challenging than ZSL. Another extended setting is Transductive Zero-Shot Learning (TZSL) [26], which relaxes the constraints of ZSL by assuming that images of unseen classes are available but labels are not available during the training phase. The setting of TZSL is too loose to alleviate the problem of data stress, and thus we do not discuss TZSL in this paper. Most studies have accommodated both ZSL and GZSL settings, and thus we collectively refer to them as Zero-Shot Object Recognition (ZSOR). Notably, several surveys on ZSOR have been presented. They comprehensively review the techniques, whereas we focus only on the element-wise studying schemes therein.
Zero-Shot Compositional Recognition (ZSCR). Another fundamental claim is that besides categorizing images, the model can also simply describe relevant states, e.g., whether an apple is ripe or not, or whether a bread is sliced or not. It is an important step for the model to truly understand the world and a key part of building human-like intelligence [27]. States are relevant properties of objects that describe the discrepancy between individuals. It means that while distinguishing between different classes, it is also necessary to recognize the characteristics of individuals in the same class. State recognition can provide more fine-grained information, which is important for model detection, perception and simulation of real environments. However, data collection is a major bottleneck. Each individual has multiple states, which means the number of samples for a complete dataset will grow geometrically. Obviously, it greatly hinders realistic applications. In contrast, humans can reason about the state of an object without having seen all the images. For example, if one has learned about sliced apples, it is easy to recognize sliced oranges. Inspired by that, Compositional Zero-Shot Learning (CZSL) [2, 3, 5] aims to explore the combinational reasoning capabilities of automated models. In the training phase, the model utilizes partial images and corresponding combinational labels for learning. In the testing phase, it needs to recognize different combinations, including seen as well as unseen combinations. Note that in the training set, every state or object has appeared at least once. Conventional CZSL is set up in such a way that the visual samples of unseen combinations are missing but possess a labeling prior, which does not correspond to real scenarios. Therefore, Open-World CZSL [28] extends the setting by expanding the labels tested to the full set, i.e., irrational label combinations, such as sliced cars, may occur, which the model needs to learn to cull out on its own. The technologies of the two are almost crossover, so we collectively refer to them as Zero-Shot Compositional Recognition (ZSCR).
Foundation model-Based Open-world Recognition (FBOR). Despite many remarkable contributions to open deployment, ZSOR and ZSCR are still insufficient to satisfy the notion of complete openness. Their objectives are fixed and the classes to be predicted are well-defined. Can the models learn the broad basics as humans do and thus adapt seamlessly to unknown tasks. The emergence of large vision-language pretraining models (VLMs) [29, 30] offers a promising solution to this problem. During training, the models focus on vision-language interactions and learns large volumes of element-wise knowledge aimlessly, just as humans observe and learn about the world. Surprisingly, the pretrained models have excellent zero-shot generalization ability, which can excel at a variety of unfamiliar image recognition tasks without further training. It has led to a wide range of applications, especially for low-shot problems [31, 32]. However, recently some studies [33, 34] point out that the zero-shot generalization ability of VLMs is not fully exploited. To this end, Foundation model-Based Open-world Recognition (FBOR) is developed to work on finding the limit of the zero-shot performance of foundation models.
Figure 2: An overview of organization.
II-B Comparison
Differences. ZSOR and ZSCR run in relatively closed environments, i.e., the categories or states to be predicted are known. Despite some studies proposing open-world ZSCR, the essence is just to extend the label set from a subset to the full set. FBOR, on the other hand, is task-agnostic, with the need to learn a wider range of knowledge ahead of time to cope with unknown challenges. In general, the element annotations of ZSOR are much larger and more detailed, while ZSCR contains only individual state and category annotations. However, the element annotations of ZSOR are class-level and thus may shift on individual instances, while the annotations of ZSCR are instance-level. In addition, the goal of ZSOR is essentially to recognize objects composed of different attributes, whereas the goal of ZSCR is to recognize combinations composed of different states and objects.
Commonalities. As mentioned earlier, element-wise representation and reasoning is motivation shared by several tasks, i.e., learning the visual meta-concepts corresponding to text annotations, and then combinational reasoning to recognize new upper-level concepts. In addition, many techniques are common across several tasks, such as attention mechanism, graph modeling, etc., which we discuss further in Section VI-C.
II-C Challenge
Fine-grained visual analysis is a critical challenge for Element-Wise Zero-Shot Image Recognition. Element-wise learning requires detailed visual anatomy to echo the text to capture the underlying meta concepts. The ideal solution is a one-to-one visual-text studying strategy, which is impractical to achieve. ZSIR has fine-grained textual information to describe the element-wise semantics. However, constrained by huge labor and time costs, images typically are lacking corresponding element-wise annotations. Therefore, text and visual regions are learned from each other in a weakly supervised manner, and their matching accuracy cannot be guaranteed. In addition, inevitably, irrelevant regions such as the background interfere with the quality of learning. Incorporating fine-grained visual analysis into ZSIR is a promising scenario, which involves visual localization, visual representation, and visual-text correlation modeling. Localization is searching those valuable regions to exclude the interference of noise, forcing the model to focus on the key parts [35, 36]. Representation is processing the contextual relationships between regions to obtain the optimal semantic embeddings [37, 38]. Visual-text correlation modeling is to establish the link between visual and text, involving the design of semantic space and mapping functions [39, 40]. The problems faced include: 1) the complexity of the elements. Compared to class-wise annotations, the number of element-wise annotations tends to be larger and the corresponding visual features are more subtle. For example, in object recognition, the number of attributes can be up to hundreds [9, 41], and they describe semantic information of different dimensions and scales. In addition, the processing of abstract attributes is also a challenge due to the difficulty of finding the corresponding explicit visual regions. 2) visual occlusion and overlap. Visual occlusion occurs between objects and backgrounds or between elements and elements, which may result in the absence of critical visual regions. Visual overlap refers to multiple semantic expressiveness of the same region [42]. For example, when recognizing red car, it is hard to completely divorce the visual information of red and car cause they can be expressed by the same visual part. 3) visual interaction. The independent semantics of visual regions are vague or even ambiguous. For example, if a piece of skin is cut off from a cow’s trunk, it is hard to recognize which part or species it comes from. Thus, the semantic meaning of regions depends to varying degrees on the whole. At the same time, there are structural and physical logical relationships between regions that make independent visual representations suboptimal.
Domain shift is another long-standing issue, which refers to the distributional discrepancy of visual features when mapped into deep semantic space [43]. The conventional paradigm of ZSIR is to map seen visual features to semantic embeddings and associate them to corresponding text annotations. When deployed to the open environment, if unknown visual samples are mapped to the vicinity of similar semantic embeddings, the corresponding text can be indexed and thus the correct label is inferred. However, modern models favor seen classes and thus learns biased mapping functions [44], making the semantic distance between unseen and seen samples too large, i.e., domain shift. It misleads the model to recognize the embeddings of unseen samples as outliers or noise and thus to generate wrong judgments, which is a huge obstacle to generalize the model to the open world. One factor contributing to domain shift is the imaging mechanism and environment [45]. Different cameras, viewpoints, lighting, backgrounds, and many other complex imaging factors may cause the same thing to appear as diverse visual characteristics. For example, the texture of an object photographed in clear weather will be brighter than in dim weather. Humans are able to counteract the style variation easily, but deep models are sensitive. To mitigate this problem, domain adaptation and domain generalization have proposed many fruitful solutions [46, 45, 47]. However, for ZSIR, it is more challenging since domain shift and unknown class prediction need to be accommodated at the same time. Another factor that leads to domain shift is even more troublesome, namely context dependency. Things expressing the same semantic meaning may present very different visual characterizations in different contexts [48, 49]. To give a few examples, the feathers of a chicken are a completely different color when young and when adult; the tail appears hairy and straight on a horse, but smooth and curly on a pig; old presents itself as a rusty patina of metal on a car, but as a fold of skin on a person. This strong contextual dependency is a major hindrance when generalizing a property from one thing to another.
II-D Organization
Fig. 2 illustrates the organization of this paper. Section III provides a comprehensive description of the techniques related to zero-shot object recognition, including the problem formulation and the introduction to the algorithms in the three main directions. Section IV provides a review of the zero-shot compositional recognition approach, also including problem formulation and the introduction to the algorithms in the three main directions. Section V provides an introduction to the background of foundation model-based open-world recognition, followed by a demonstration of three VLM-related zero-shot tasks and recent progress. In section VI, we show the technical details, including commonly used datasets, technical details, and discuss the techniques they share. Section VII shows the related application scenarios. Finally, we share our views on the future directions in section VIII.
III Zero-Shot Object Recognition
III-A Problem Formulation
Given the seen domain $\mathcal{D}^{s}=\{(x^{s},y^{s},a^{s})|x^{s}\in\mathcal{X}^{s},y^{s}\in\mathcal{Y}^{s},a^{s}\in\mathcal{A}^{s}\}$, where $\mathcal{X}^{s}$, $\mathcal{Y}^{s}$, and $\mathcal{A}^{s}$ denote visual samples, category labels, and semantics (e.g., a set of attributes), and similarly, let $\mathcal{D}^{u}=\{(x^{u},y^{u},a^{u})|x^{u}\in\mathcal{X}^{u},y^{u}\in\mathcal{Y}^{u},a^{u}\in\mathcal{A}^{u}\}$ denote the unseen domain. Notably, $a^{s}\cup a^{u}\in\mathbb{R}^{m\times k}$ are class-level semantics annotations, where $m$ denotes the number of categories and $k$ is the number of attributes. Without loss of generality, the task of ZSOR can be modeled as learning a mapping/relational function $\Psi:\mathcal{X}\to\mathcal{A}$, wherein $\mathcal{X}^{u}$ is strictly inaccessible for training. During inference, the learned function $\Psi$ is applied to recognize samples from the unseen domain only, i.e., ZSL, or from the joint of both seen and unseen domains, i.e., Generalized ZSL (GZSL).
The success of ZSOR relies on the sharing property between $\mathcal{A}^{s}$ and $\mathcal{A}^{u}$, which act as the bridge from seen to unseen domains. An intuitive idea is to learn the visual representations corresponding to attributes, which is the notion of element-wise learning. However, due to the single target and relatively closed environment of object recognition, learning for $\Psi$ can be transformed into a plain class-wise distribution matching problem. Class-wise distribution matching concerns neither the discrepancies of the bottom attributes nor the variations of fine-grained vision. It treats both visual and attribute features as a whole, and the objective is to reduce the distributional distance between the two so as to learn an unbiased mapping function. Mainstream techniques include semantic space design, generative methods and regularization. The goal of semantic space design is to search for a suitable low-dimensional space that perfectly matches visual and attribute embeddings. Common practices are one-way mapping [50, 51], which projects visual/attribute into the space of the other, and common space [44], which projects both visual and attribute into a common space. The goal of the generative approach is to simulate the visual features of unseen classes [52, 53]. The basic approach is to train a generator that produces visual features based on attributes with the help of networks such as GANs [54] and VAEs [55]. The goal of regularization is imposing constraints to improve the mapping performance or robustness of the model. A popular approach is to design a specific loss function. For example, Liu et al. [56] uses an information entropy loss for the unseen class thereby reducing uncertainty.
Despite the impressive results achieved by these methods, they provide limited insights for micro-operations and bottom-up analyses, and are therefore out of the scope of this paper. We recommend the reader to refer to other surveys [24, 6] for a deeper understanding of the above methods.
Depending on who the method operates on, we categorize the element-wise techniques into three broad classes, including Visual Component Analysis, Visual-Attribute Matching, and Attribute Component Analysis. Then the methods are further categorized into different subclasses based on the specific techniques they take.
III-B Visual Component Analysis
Visual component analysis is the refined dissection of an image or its features, involving localization, modeling, and so on. Images often contain a large amount of redundant information, only a few of which are decisive for discrimination. At the same time, images lack fine-grained annotations. Therefore, the goal of visual component analysis is to reduce the interference of irrelevant information as well as noise, and prompt the model to focus on those critical visual parts. Commonly used methods can be classified into the following three categories: local visual attention, data manipulation and graph modeling.
III-B1 Local Visual Attention
Figure 3: A schematic diagram of local visual attention. Visual features are passed through a sub-network to generate multiple masks, which are then multiplied to obtain enhanced features.
People selectively focus on the most attractive parts of things when they look at them, and these places usually contain the most discriminative information. It is this behavior that the attention mechanism mimics to make the model concentrate on the most significant areas, thus improving the efficiency and accuracy of information processing. Attention mechanisms have a wide range of applications in computer vision, natural language processing and other fields. The goal of local visual attention is to utilize the attention mechanism to identify the most valuable visual regions and reduce the influence of background and irrelevant regions. Specifically, local visual attention generates multiple masks for visual features. Each mask is a normalized weight matrix that reveals a significant region. Enhanced features can be obtained by weighting the original features using mask, thus improving the discriminability (Fig. 3). Suppose that $x\in\mathbb{R}^{C\times H\times W}$ denotes the visual feature, where $C$, $H$, and $W$ are the channel, height, and weight, respectively. Let $M=\mathcal{F}(x)$ represents the attention mask, where $\mathcal{F}$ denotes a small learnable network, e.g., $1\times 1$ convolutional network, and $M\in\mathbb{R}^{N\times H\times W}$ denotes $N$ attention masks. Afterward, multiplying the masks with the original features yields $N$ regional features, which can be expressed as:
$x_{region}=\{xm_{1},xm_{2},...,xm_{N}\},$ (1)
where $x_{region}\in\mathbb{R}^{N\times C\times H\times W}$ and $[m_{1},m_{2},...,m_{N}]=M,m_{i}\in\mathbb{R}^{H\times W}$. How to use these regional features to improve model’s performance. AREN [36] considers certain mask learned regions to be unimportant and therefore adopts an adaptive threshold to further filter out unimportant information, i.e., region features smaller than the threshold are set to 0. RSAN [57] expects the regions learned by each mask to correspond to the attributes one by one. It uses max-pooling to get the representative value in each region feature and concatenates them into a one-dimensional vector, after which it uses Mean Squared Error Loss to minimize the distance between this vector and the attribute embedding. Considering the existence of topological relationships between regions, RGEN [37] introduces graph modeling to maintain their potential connections.
III-B2 Data Manipulation
The objective of data manipulation is the same as local visual attention, i.e., to search and localize the most critical regions in an image. The difference is that local visual attention is weighted at the semantic feature level, whereas data manipulation operates at the image level. An intuitive scheme is to recognize parts of an object, such as the head, body and claws of a bird, with the help of detection techniques. The detection network can provide coordinates of multiple parts simultaneously. Afterwards, image patches are cropped based on the coordinates or the relevant locations are localized in deep features. These parts introduce a large amount of fine-grained semantic information that assists in facilitating the classification ability of the model  [58, 59]. Additionally, the technique also improves the performance of generative methods, which struggle to capture fine-grained visual information differences. For example, AGAA [60] utilizes a detector to extract region information that is then sent to a discriminator for training along with class embeddings generated by a generator. The application of the detection technique is very limited because the performance of the detector relies on part-level visual annotations, which is undesirable in most scenarios.
To alleviate this problem, a series of unsupervised detection methods have been proposed. LDF [61] employs a coordinate generating network that generates coordinates directly from image features. Specifically, suppose $x\in\mathbb{R}^{C\times H\times W}$ is the visual feature, and the coordinate can be expressed as:
$Z=[z_{h},z_{w},z_{l}]=\mathcal{F}(x),$ (2)
where $\mathcal{F}$ is a learnable network. $Z$ is the window, $z_{h},z_{w}$ denote the coordinates, and $z_{l}$ denotes the length of the region. After that, the regions located at the coordinates are zoomed in and re-fed to the classification network. SGMA [35], on the other hand, unites local visual attention techniques to generate multiple region coordinates at once. It first generates multiple attention masks and then uses the attention masks as inputs to get the corresponding coordinates. Finally, the original image is cropped according to the coordinates. Furthermore, SR2E [62] and ERPCNet [63] introduce the reinforcement learning paradigm, which induce the cropping behavior of the model by setting up special rewards and penalties to obtain the most discriminative sub-image. POPRNet [64] takes unsupervised detection technique. It generates multiple proposal regions for an image and then sorts and selects Top-K region features based on category confidence. ZSLViT [65] utilizes the patch split mechanism of ViT [66] to get multiple region features and later filters the key visual regions by calculating the attention weights between the regional and class tokens.
Unlike the above approaches, to capture the key visual semantic embeddings, VGSE-SMO [67] takes a patch clustering approach. It utilizes an unsupervised compact watershed segmentation algorithm [68] to decompose images into multiple patches, after which individual patches are linked to class labels via a classification network. By doing so, the model implicitly learns the relationship between classes and fine-grained visuals.
III-B3 Graph Modeling
Visual region localization promotes a model that focuses on important local features, however, fails to take into account potential relationships between regions. For example, a bird has biological topological connections between its various body parts. Independent regional representation loses substantial semantic information by narrowing the field of view, which may lead to semantic ambiguity or disambiguation. Sharing information between regions can be a useful method to correct semantic information by compensating for the loss of field of view. Fortunately, graph modeling techniques is able to well handle such irregular non-geometric data with a structured message propagation mechanism [7]. In recent years, with the birth and development of graph convolutional networks (GCNs) [69], graph modeling techniques have been widely used in various computer vision tasks. Graph convolutional networks ingeniously integrate the concept of graph into convolutional neural networks. Suppose $W^{(l)}$ denotes the parameters of the $l$-th layer of GCNs, the output of the $(l+1)$-layer can be expressed as:
$H^{(l+1)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}H^{(l)}W^{(l)}),$ (3)
where $A$ is the adjacent matrix and $\tilde{D}$ denotes its degree matrix. $\sigma$ denotes the activation function, and $H^{(l)}$ is the output of the $l$-layer of GCNs. It can be seen that the nodes share information through an inside-out information aggregation mechanism through the graph network.
Regional visual features are natural graph nodes and thus can be well embedded in graph modeling techniques. For example, GKU [38] constructs a graph for each sample to capture fine-grained local structural relationships. They redefine the visual-semantic mapping problem as a graph-semantic mapping problem, thus suppressing the influence of partial visual domain shift. However, GKU requires the dataset to be supported by fine-grained regional annotations. Without the supervision of manual annotations, RIAE [70] incorporates all regions into the graph node set. To mine the relationships between regions, it introduces cosine similarity to measure the weight of an edge between any two region nodes. GNDAN [71], on the other hand, first locates the visual regions corresponding to the attributes with cross attention techniques, and then later uses the graph for modeling.
III-C Visual-Attribute Matching
The objective of visual-attribute matching is to find appropriate visual features for each attribute, and to learn fine-grained element-wise representations by searching and matching pairs of visual-attribute embeddings. Compared to visual component analysis, visual-attribute matching has a more explicit motivation to interact with attributes. At the same time, visual-attribute matching techniques have a stronger perception of subtle variations between attributes, leading to better generalization to unseen domains. Such text-induced visual element learning is more compatible with human learning habits and more challenging. Existing approaches can be categorized into cross attention and prototype learning.
III-C1 Cross Attention
Figure 4: A schematic diagram of cross attention. Attribute embeddings and region features compute similarity to obtain multiple attention maps.
Attention mechanisms have been used in visual component analysis to localize discriminative regions that are friendly to class labels but fail to capture subtle visual differences between attributes. Therefore, cross attention is introduced to discover visual regions of interest for attributes. Cross attention allows for simultaneous processing of different sources of information and efficient modeling of their dependencies. Among them, dot product attention, due to its high computational efficiency and low space cost is widely applied in many advanced models such as Transformer [72]. The core of dot product attention is to compute the similarity between query and key-value pairs to assign attention weights, and the similarity measure is the dot product of vectors. Given a query series $Q$, a key series $K$ and a value series $V$, the attention score can be formulated as:
$Attention(Q,K,V)=softmax(\frac{QK^{\mathsf{T}}}{\sqrt{d_{k}}})V,$ (4)
where $d_{k}$ denotes the dimension of the key vector which is used to scale the dot product result. When $Q$, $K$, and $V$ all come from the same input series, the attention score learns the dependencies between the parts of the series, i.e., self-attention. And when there are two different input series involved in the computation, the attention score learns the dependencies between the different series, i.e., cross attention.
In order to uncover the interactions between attributes and visual regions, the cross attention technique is naturally introduced into ZSOR (Fig. 4). For example, DAZLE [39] proposes a dense attention technique that assigns weights to each region based on attributes. Specifically, it utilizes the features output from the last layer of the convolutional network and treats each point on the feature as a region. The similarity between attributes and regions is then computed with cross attention. Suppose $x\in\mathbb{R}^{C\times r}$ denotes the visual feature with $r=H\times W$ regions. Let $a\in\mathbb{R}^{d\times m}$ denote the attribute vector, where $m$ is the number of attributes and $d$ is the vector dimension. The cross attention score is computed as:
$Attention(a,x)=\frac{\exp(a^{\mathsf{T}}wx)}{\sum^{r}\exp(a^{\mathsf{T}}wx)}x,$ (5)
where $w\in\mathbb{R}^{d\times C}$ is a learnable matrix which ensures the visual and attribute vectors are in the same dimension space. The weighted visual features reflect attribute preferences. After that DAZLE uses the attention scores of all the attributes to classification. GEM [73] employs the same approach, however it retains the maximum attention scores with max-pooling during classification. Composer [74] synthesizes the features of unseen classes with critical regions identified by attention mechanism, while AREES [75] exploits the regional features to improve the performance of VAE.
Some studies resort to ViT [66] to model dependencies between vision and attributes [76, 77, 78]. The essential cross attention mechanism is identical. The difference is that ViT divides the image into patches, each of which is considered as a visual region.
In addition, bidirectional attention is explored in depth. It is important to know not only which region the attribute is attending to, but also which attribute the region is attending to. Based on this idea, several methods are developed. For example, MSDN [79] proposes a mutual distillation network to align attribute-region attention. The attribute-to-region attention algorithm is the same as Eq. 5, and the region-to-attention to attributes is formulated as:
$Attention(x,a)=\frac{\exp(x^{\mathsf{T}}wa)}{\sum^{m}\exp(x^{\mathsf{T}}wa)}a,$ (6)
where $w\in\mathbb{R}^{C\times d}$ is a learnable matrix. Afterwards, it uses two attention scores to compute the respective classification results and aligns them with distillation loss. Ge et al. [48] also integrate bidirectional attention techniques to improve the visual-attribute matching ability of model.
III-C2 Prototype Learning
The objective of prototype learning is to store representative samples and perform tasks such as classification and regression by calculating distances to them [80]. A prototype is a standardized sample or example that represents a category or type. For example, the mean vector of a collection of vectors or the center of mass of a distribution is a prototype. Prototype learning imitates the human mindset of analogical reasoning, i.e., reasoning and summarizing by analogy with typical examples. Its main advantage is its stability in dealing with complex data distributions, especially when there is overlap or imbalance between data categories. Therefore, for tasks such as out-of-distribution detection [81] and low-shot learning [82], prototype learning is considered to have stronger robustness.
For visual-attribute matching, there is semantic overlap in the visual region. The same region may contribute to multiple attributes. Although cross attention is able to measure the contribution value by similarity, the lack of an explicit means of visual decoupling may lead to pseudo-correlation of visual-attributes. Therefore, learning a separate visual prototype for each attribute is a promising solution. When the prototype is acquired, the visual region can decide which attribute it belongs to by calculating the distance between the region vector and the prototype. It can be observed that cross attention discovers pairing relations through the similarity between attribute embeddings and visual embeddings, while prototype learning discovers visual-attribute pairing relations indirectly through the similarity between visual embeddings and visual prototype embeddings.
APN [83] first introduces prototype learning to visual-attribute matching. It initializes a learnable prototype vector for each attribute and then calculates the similarity between the visual region and the prototype. After that, the largest region vector is taken as the attribute vector, and then regresses with the attribute embedding. In addition, it incorporates a decorrelation loss [84] to ensure the independence between individual attributes. DPPN [85] considers that the prototype possesses a seen-domain preference, which cannot capture the delicate instance-level attribute shift. Therefore, it is necessary to equip each instance with exclusive attribute prototypes. It proposes a progressive attribute localization method, which dynamically adjusts the prototype by locating attributes through the prototype while updating the prototype in turn. In addition,  [48, 86, 87, 88] also achieve favorable results by using prototype learning.
III-D Attribute Component Analysis
Attributes are the only bridge between seen and unseen classes, and they are also important supervisory signals for learning fine-grained visual representations. Therefore, the quality of the attribute space heavily determines the upper limit of the performance of a recognition system. However, most of the attributes are usually created manually or pretrained without special adjustment or filtering. Attributes contain a lot of redundant information and noise, and the complex relationships between attributes are not sorted out. As a result, the learned attribute space is not optimal. Attribute component analysis attempts to improve the above situation by adjusting or designing the attribute space. The main approaches include attribute enhancement and attribute refinement.
III-D1 Attribute Enhancement
The purpose of attribute enhancement is to optimize attribute embeddings to boost the discriminability of attributes, and the main techniques include key attribute localization and attribute space optimization. The motivation of key attribute localization is similar to that of visual attention, which aids in screening by identifying and localizing the most critical parts. The cross attention mechanism mentioned in Equation 6 is one of the attribute localization techniques. In addition, LFGAA [89] designs a latent guided attention module. It assumes that each image has different attention for attributes and therefore introduces multiple subnetworks to generate exclusive attribute weights for each image. Attribute space optimization is to adjust the pretrained attribute embeddings. The original attribute embeddings have semantic overlaps affecting the discrimination, especially for fine-grained datasets. For example, a large number of attribute features are shared between different bird species, which only have minor visual variations in some of the attributes. And the original attribute embeddings are difficult to capture such attribute differences resulting in a sub-optimal space. For this reason, AREES [75] delineates the attribute embeddings with similar semantic distances into a cluster by K-means clustering algorithm, and the attributes in each cluster are subtracted from the cluster’s center of mass, i.e., the mean vector, to obtain the enhanced attribute vector. APNet [90] introduces graph modeling to discover the logical relationships between classes, and updates the attribute vectors by propagating the attribute features of neighboring nodes.
III-D2 Attribute Refinement
Most algorithms just learn attribute annotations as normal labels or optimize the attribute embedding space. However, the potential irrationality of attribute annotations has not been thoroughly investigated. Since most attributes are annotated by humans, the gain of description from human perspective for deep learning is unknowable. Therefore, attribute refinement attempts to provide insights from the perspective of attribute annotation design. However, there are fewer related studies. MCZSL [91] points out that the quality of manual annotations is uncontrollable and time-consuming. Therefore, it attempts to design an automated attribute annotation scheme. Specifically, it obtains attribute descriptions from online media such as Wikipedia and designs a series of filtering strategies to filter out inappropriate attributes. DEDN [92] proposes a divide-and-conquer attribute learning scheme. It argues that attribute annotations are confusing and lead to semantic asymmetry problems. For example, for the attributes describing scenes, {studying, swimming} and {stone, tree} are clearly not in the same dimension, with the former denoting function and the latter denoting object. Therefore, it designs a dual expert network in which one expert network focuses on processing attribute information in different dimensions.
IV Zero-Shot Compositional Recognition
IV-A Problem Formulation
Suppose $\mathcal{S}=\{s_{1},s_{2},...,s_{n}\}$ is a set containing $n$ state elements and $\mathcal{O}=\{o_{1},o_{2},...,o_{m}\}$ is a set containing $m$ object elements. Then a compositional set can be obtained as $\mathcal{Y}=\{(s_{1},o_{1}),(s_{1},o_{2}),...,(s_{n},o_{m})|s\in\mathcal{S},o\in\mathcal{O}\}$, which containing $n\times m$ compositions. Now we have a seen domain $D^{s}=\{(x^{s},y^{s})|x^{s}\in\mathcal{X}^{s},y^{s}\in\mathcal{Y}^{s}\}$, where $\mathcal{X}^{s}$ denotes the training image set and $\mathcal{Y}^{s}\subseteq\mathcal{Y}$ is the label set. Similarly, we have an unseen domain denoted by $D^{u}=\{(x^{u},y^{u})|x^{u}\in\mathcal{X}^{u},y^{u}\in\mathcal{Y}^{u}\}$, where $\mathcal{X}^{u}$ denotes the test image set and $\mathcal{Y}^{u}\subseteq\mathcal{Y}$ is the label set. Data in the seen and unseen domains are disjoint, i.e., $X^{s}\cap X^{u}=\phi$ and $Y^{s}\cap Y^{u}=\phi$. The objective of ZSCR is to generalize to the unseen domain by learning the element-wise knowledge in the seen domain. Specifically, for closed-world compositional recognition, the goal is to learn a mapping function $\Psi:\mathcal{X}\to\mathcal{Y}^{s}\cup\mathcal{Y}^{u}$, wherein $\mathcal{X}^{u}$ is strictly inaccessible for training. Typically, the size of $\mathcal{Y}^{s}\cup\mathcal{Y}^{u}$ is smaller than that of $\mathcal{Y}$ since some compositional are not feasible. For open-world compositional recognition [28], the goal is to learn a mapping function $\Psi:\mathcal{X}\to\mathcal{Y}$, assuming that which compositions are infeasible is unknown.
We divide the cutting-edge techniques of ZSCR into three categories based on their core components, including Decoupling Representation, Dependency Modeling, and Relationship Modeling. We then further split the subcategories based on specific technical details or targets of action.
IV-B Decoupling Representation
A common baseline approach is to map the image to a deep semantic space and then compute the similarity to the text embeddings and optimize with cross entropy loss. Briefly, the goal is to learn the posteriori probability $p(y|\psi(x))$, where $y=(s,o)$ and $\psi$ is a function mapping the image into the semantic space. This approach takes into account the mapping relation of compositional embeddings, yet lacks the ability to deconstruct and reorganize the visual features, thus making it difficult to generalize to the unseen domain. In other words, during training phase, the model will seek the simplest path for optimization, i.e., learning the correspondence between hybrid visual features and labels, without considering the independent effects of states and objects. Thus, the model tends to capture correlations between states and objects that are preferred for the seen domain. For example, if the model has only seen sliced apple, it may tightly associate sliced with apple and fail to recognize sliced orange. An intuitive idea is to predict the labels of states and objects separately, excluding semantic interference with each other, which is the goal of decoupling representation. Decoupling representations strive to untangle states and objects, encouraging the model to extract semantic-independent visual representations. By recognizing states and objects individually, all seen or unseen compositions are in turn recognized. Decoupling representation is to learn the posteriori probabilities $p(s|\psi_{s}(x))$ and $p(o|\psi_{o}(x))$, where $\psi_{s}$ and $\psi_{o}$ are mapping functions. Mainstream approaches can be categorized as network design and attention separation.
IV-B1 Network Design
Many researches resort to modifying the network structure to decompose visual features. Nan et al. [93] add two sub-networks at the bottom of the backbone network to extract state and object features, respectively, and then each predicts the corresponding labels and is trained with cross-entropy loss. CASUAL [42] employs a similar network structure. The difference is that it introduces a text-to-image mapping network that reverse maps text embeddings into visual space. Meanwhile, it uses the Hilbert-Schmidt Information Criterion, which is a non-parametric method for estimating the statistical dependence between samples of two random variables, to optimize training for ensuring state and object independence. ProtoProp [94] uses a prototypical network to learn independent representations. Specifically, regional features of images and prototypes of objects and states are computed for similarity and later optimized with cross-entropy loss. Hu et al. [95] propose a two-layer decoupling network. First, three sub-networks are used to extract state, object and composition features respectively. In the second layer, the state and object features are fed into a reorganization network to synthesize into a composition, while the composition feature is fed into two other sub-networks for further decoupling. CAILA [96] uses CLIP as a visual coder, which inserts three adapters in the middle of the network to capture the state, object, and composition semantic information, respectively. It is worth mentioning that the three-branch network has become a common infrastructure for most approaches, i.e., three sub-networks extract state, object, and composition visual features, respectively [97].
IV-B2 Attention Separation
The motivation for attention separation is to dissect semantic features by capturing the parts of objects and states that are attended to separately. OADis [98] introduces a cross-sample attention mechanism to refine features. It equips each sample with two companion samples, one with the same object and different state, and the other with the same state and different object. Afterwards, it performs cross attention with the corresponding samples separately to refine the semantic features. ADE [99] adopts a similar training paradigm. Unlike pre-matched pairs, Jing et al. [100] utilize image retrieval techniques to find suitable matching samples for object and state branches, after which they are used to enhance the representations. Zhang et al. [101] propose a gradient-based attention decomposition method. They view compositional recognition as a domain generalization problem, and one of the solutions is to learn domain-invariant features. To this end, they measure the variation of features by evaluating the difference of the back-propagation gradients of sample pairs to recognize the domain-invariant part. DFSP [102] firstly decouples text embeddings by prototype computation, and after that, uses the text embeddings and image features for cross-attention in order to decompose the visual features. Similarly, Troika fuses multi-path prompt and cross attention to optimize text embeddings.
IV-C Dependency Modeling
Decoupling techniques pursue the pure semantic representations, while mitigating the bias toward seen compositional concepts, introduce the bias toward single concept, i.e., state or object. This bias comes from the fact that the same concept is represented differently in diverse contexts. Due to the existence of complex acting mechanisms in the real world such as physical systems and biochemical reactions, the visual presentation of the same state on various objects varies greatly. For example, old is represented as metal rust spots on car, while it is shown as skin wrinkles on person. Obviously, concise text annotations cannot capture such subtle differences. And the independent visual representations learned by the model from the seen domain are biased, hindering the recognition accuracy on the unseen domain. To address this problem, many studies devote to modeling the dependency relationship between states and objects. By introducing a priori information about states or objects to correct each other’s semantic representations or classification confidence, they can be better generalized to the unseen domain. In short, the goal of dependency modeling is to learn the conditional probabilities $p(s|\psi_{s}(x),o)$ and $p(o|\psi_{o}(x),s)$. The relevant technologies according to the implementation objectives can be categorized into feature optimization and aid classification.
IV-C1 Feature Optimization
Figure 5: A schematic diagram of dependency modeling [103]. Object information is used as a signal to guide the extraction process of state features.
The goal of feature optimization is to adjust each other’s feature representations by leveraging the semantic relevance of states and objects. Decoupled state/object features are expected to capture dependent semantic details while maintaining their own discriminative semantic information. Visual feature optimization is to correct the visual embedding distribution to avoid overfitting the seen domain caused by the model focusing on the single visual feature. For example, Panda et al. [104] set up multiple encoders between the state and object branching networks with the aim of transmitting object features to the state network in order to model the dependency of the state on the object. The same idea is taken by CoT [103], which uses object information as cueing signals while introducing an attention-based method to capture state-related visual features (Fig. 5). Another approach is to optimize text features. Since text annotations are usually pre-extracted, they do not have the ability to perceive dependency relationships. To this end, Nagarajan et al. [8] model states as operators, i.e., encode the state as a mapping function rather than an embedding, and thus explicitly learn the semantic embedding of an object ”after it has been operated on”. SymNet [105] introduces two sub-networks, the decoupling network and the coupling network, which are used to learn independent and compositional representations of objects and states. CANet [106] recognizes the object labels of an image firstly and then corrects the text embedding of the state by using the object text embedding as a priori condition.
IV-C2 Aid Classification
Compared to adjusting the feature distribution, aid classification aims to directly introduce semantic information of objects/states to help the classifier to make discriminations. In earlier work, Chen et al. [2] utilize support vector machines (SVMs) for compositional recognition. They learn the weights of SVMs by capturing state dependencies on objects and generalize to unseen compositions. SAD-SP [107] takes an intuitive approach by introducing an attention module in addition to the conventional decoupling network. State features are used to predict object soft labels and vice versa. The predicted soft labels are then used to weight the predicted output of the conventional network. ProCC [108] performs information exchange at each layer of the state/object classifier. Specifically, it passes the output of one layer of the state classifier through a one-dimensional convolution and softmax, and then the output of the corresponding layer of the object classifier is multiplied with it and used as the input of the next layer. And vice versa.
IV-D Relationship Modeling
Dependency modeling promotes the generalizability of models by exploring the intrinsic one-to-one state-object connections. Naturally, we wonder whether many-to-many complex mesh relationships, i.e., interactions between multiple objects or states, contribute to model performance. Such relationships describe the latent logical relations of elements in the real world due to various factors, such as subject affiliation, visual similarity, co-occurrence, etc. To name a few, we would spontaneously categorize broccoli and spinach, separating beef since they belong to different subjects; a fork and a spoon are more visually similar than an umbrella, and also they have a higher probability of co-occurrence. Intuitively, an understanding of this relations facilitates logical inference. It is because humans have a remarkable ability to draw analogies, make connections, and imagine, thus reasoning and recognizing by tapping into the relationships between different elements. When we see a furry black bear and know the similarities between polar bears and black bears, it is easy to migrate the state of furry to polar bears. Driven by this idea, relationship modeling aims to enhance the inference and representation capabilities of models by uncovering and maintaining potential relationships between states and objects. The dominant techniques can be summarized as cross attention, graph modeling and contrastive embedding.
IV-D1 Cross Attention
Cross attention is used to capture similarities between text embeddings and thus correct text representations. BMPNet [109] introduces the cross attention mechanism to improve independent object/state representations. Specifically, for a state, the attention weights of the state and other states are first computed, and then the attention weights of and object are computed, resulting in an enhanced state representation. The object ditto. Unlike object/state relational modeling, CAPE [110] aims to reveal co-occurrence relations between compositions. It adapts the compositional features by calculating the attention weights between the text embeddings of the compositions.
IV-D2 Graph Modeling
As mentioned in the previous section, graphs are excellent at working with structured data and thus can be well adapted to the relationship modeling task. In earlier work, Cruz et al. [16] develop a logic tree to model relations between states and objects. They want the model to learn the notion of connectives such as {and, or, not} to understand the potential connections between nouns. CGE [111] models the entire set of objects and the set of states. It treats each object and state as a graph node and then identifies edges between nodes by the composition labels. Graph convolutional network is later introduced to propagate the information. CVGAE [112] expects the model to be able to determine on its own whether there is a connection between objects and states, and thus be adaptable to the open-world compositional recognition task. It designs a reconstruction task where the reconstruction network recovers the adjacency matrix based on latent variables. GIPCOL [113], on the other hand, fuses prompt learning and graph modeling. It employs CLIP as backbone and takes the graph-encoded text embeddings as prompts to train. The graph neural network is involved in the learning process thereby optimizing the text embeddings to better fit the task.
IV-D3 Contrastive Embedding
Figure 6: A schematic diagram of contrastive embedding [49].
Contrastive embedding utilizes contrastive learning to optimize the semantic space and implicitly learn the relationships between samples. Contrastive learning [114], as a representative unsupervised learning method, demonstrates powerful semantic representational ability. It aims at clustering similar samples together and separating different samples. The learning mode is to constrain the semantic distance by comparing positive and negative examples, and the learned semantic distance reflects the affinity between the samples, i.e., the semantic relationship. Therefore, many areas exploit contrast learning to optimize semantic embeddings. Unlike the attention mechanism and graph modeling described above, contrast learning does not pre-define the relationship between samples explicitly, but rather iteratively adjusts the semantic distances between samples to place them in appropriate positions. The performance of contrast learning greatly depends on the number and quality of negative examples.
Wei et al. [49] propose a quintuplet loss to optimize the text embedding space. Specifically, they use a sub-network to synthesize state and object texts into compositions. After that, for an image, its label is used as a positive example and supplemented with three negative examples for contrastive learning. The negative examples are text compositions with the same state but different objects, the same object but different states, and different object and state with the target image (Fig. 6). SCEN [115] performs contrastive learning in the visual semantic space. It defines two databases to sample the negative examples, one with the same state as the target image but with different objects, and the other with the same object but with different states. After that it implements contrastive learning in the decoupled state/object embedding space respectively. Yang et al. [116] adopt the similar approach for exploring the relations among states or objects. Meanwhile, they introduce instance-level contrast learning to optimize the compositional embedding space.
IV-E Others
In addition to the above mainstream techniques, many researches provide unique perspectives to address the problem of compositional recognition. TMN [5] presents a gated network that recognizes different compositions by injecting state and object information into the network to determine the switching of pathways. Similarly, TAFE-Net [117] proposes a task-aware learning paradigm that guides the studying process by using task information as cues. Some studies are devoting to analyzing the imbalance issue. MUST [118] argues that the difficulty of recognizing elements varies, e.g., mashed banana is hard to identified as banana but relatively easy to identified as mashed. For this reason, it introduces a weighted loss to balance the effects of different elements. ProLT [119], on the other hand, points out that the data imbalance problem caused by different sample size of elements affects the recognition accuracy. To this end, it similarly proposes a novel weighted loss to solve this problem. CDS-CZSL [120] argues that for an object, the importance of different states is different. For example, sliced is more important than red for strawberry, cause red is a more general state that can be reflected in a wide range of categories. In this regard, it firstly highlights those important states by clustering and contrasting. And then a specificity-refined primitive loss is implemented to optimize the network. Moreover, several studies develop new tasks to address larger challenges [121, 122].
V Foundation model-Based Open-world Recognition
V-A Problem Description
Traditional deep vision models have achieved great results on small closed-form datasets. At the same time, large language models have seen unprecedented growth. There has been a desire to bridge the visual and language barriers and empower models to process both modal data simultaneously. Large pre-trained visual language models (VLMs) offer a promising solution [123]. The motivation is intuitive and succinct, i.e., to learn the matching relations of visual-language in a super-large image-text pair dataset to understand the element-wise knowledge of images. For example, CLIP [29] employs contrastive learning to embed images and text into the same semantic space. Surprisingly, the knowledge learned by VLMs is extremely versatile and transferable, which can be directly applied to the open-world recognition tasks. It is only necessary to associate the text description with the class name of the downstream recognition task, e.g., a photo of a [class], to achieve the purpose of classification and yield excellent results. However, the capabilities of VLMs have not been fully exploited. In recent years, there has been a proliferation of research addressing the generalization capabilities of VLMs in the open world. The mainstream research directions are as follows.
Better VLMs. Upgrading for VLMs has largely been locked into the exploration of pre-training approaches. The motivation for most studies stems from two aspects, data-efficient pre-training [30, 124, 125] and learning-efficient pre-training [126, 127, 128, 129]. The purpose of data-efficient pre-training is to explore efficient training approaches that reduce the data volume. VLMs typically require the support of large-scale datasets, e.g., CLIP trains with 400 million pairs of image-text, which is obviously a time-consuming and labor-intensive project. For this reason, some research has been devoted to developing better learning approaches that maintain excellent representation capabilities while greatly reducing the data volume. The purpose of learning-efficient pre-training can be summarized as learning better representation. Contrastive learning provides a simple yet efficient paradigm for the training of VLMs. However, whether the learned representation is optimal is a long-standing question. Researchers have attempted to parse and tackle it from different perspectives. One is to explore different learning approaches, such as reconstruction [127], generation [128]. Another is to improve the representation space from the perspective of data interaction, e.g., fine-grained image-text interaction [126], data debiasing [129].
Few-Shot Adaptation and OOD Detection. VLMs are widely used in various downstream tasks because of their excellent representation capabilities. Few-shot adaptation refers to the use of VLMs for few-shot recognition tasks [31, 32]. Specifically, by fine-tuning pre-trained VLMs with a small amount of labeled data, a specialized model that performs well on that task can be obtained. The goal of OOD detection, i.e., out-of-distribution detection, is to identify samples that are not in the target categories [130, 131]. Pre-trained VLMs are first fine-tuned on a specific recognition task. In the testing phase, the model needs to identify foreign samples that do not belong to any of the categories and assign them to the unknown class. It is important to note that the OOD does not need to recognize the specific category of the foreign samples.
Zero-Shot Adaptation. Both few-shot adaptation and OOD detection require labeled data as support to fine-tune VLMs, and zero-shot adaptation releases this limitation. Zero-shot adaptation consists of three basic tasks that apply VLMs to open environment recognition tasks under different constraints.
Zero-shot prediction: With only the category names of the task known and without any image data applying VLMs to the downstream task.
Test-Time Adaptation: Know the category names of the task and allow adjustments at the online testing phase.
Unsupervised Adaptation: Know the category names of the task and have part or entire of the unlabeled test data in advance.
In this paper, we will detail the current progress of zero-shot adaptation according to topic.
Figure 7: A schematic diagram of test-time prompt learning [132]. Adjust the prompt by minimizing the entropy of the output.
V-B Zero-Shot Prediction
VLMs have powerful zero-shot prediction capabilities. For a specific task, simply editing the category names into the prompt, e.g., a photo of [class], enables direct classification without fine-tuning. However, there is a performance gap between such a deployment and a model that is fully supervised trained on the downstream task. To address this issue, zero-shot prediction presents promising solutions. A prevalent solution is to design better prompt. The commonly used a photo of [class], while intuitive, is missing a lot of detailed information. For this reason, Menon et al. [33] propose to use a large language model such as GPT-3 [133] to generate multiple word descriptions for classes. In the testing time, the similarity between the image sample and each word description is calculated and the mean value is taken as the discriminative basis. CuPL [134] has a similar idea, which uses this description directly as the prompt for testing. CHiLS [135] also adopts a large language model, which generates multiple subclasses for each category and uses their combined similarity as the discriminative basis. Unlike the design of prompt, SuS-X [136] uses Diffusion [137] to pre-generate multiple pseudo-visual samples for each class, which serves as a support set to assist the judgment at the testing time. CALIP [34] takes the perspective of embedding space, and it introduces the attention mechanism to capture the correlation between text embeddings and visual embeddings, which effectively improves the zero-shot generalization capability.
V-C Test-Time Adaptation
The purpose of Test-Time Adaptation (TTA) is to dynamically adjust the model to progressively adapt to the downstream task during the testing process. The mainstream schemes are categorized into prompt adaptation and cache bank. Prompt adaptation is a technique that combines prompt learning and TTA. During testing, the prompt is set to learnable parameters and the parameters of the model are fixed. By continuously optimizing the prompt representation and thus adapting to the target task. TPT [132] first presents this concept by augmenting the test image into multiple views and then filtering out those with low confidence. The confidence outputs of the remaining augmented samples are averaged as the final prediction and, at the same time, the information entropy is minimized in order to train the prompt (Fig. 7). This paradigm provides a basic template for subsequent research on prompt adaptation. DiffTPT [138] argues that the augmentation approach of TPT lacks diversity, and therefore introduces the diffusion model to generate richer samples for the test image. PromptAlign [139] argues that adjusting the prompt should be done while aligning the distribution of the target and pre-training data. Therefore it introduces additional loss to align the statistics of target and pre-training data, i.e., mean and variance. SwapPrompt [140] and DART [141] use the exponential moving average algorithm to update prompt.
The goal of the Cache bank is to store the tested samples to improve the performance of subsequent tests. An intuitive approach is to cache the previous high-confidence samples as the support set to be queried by subsequent samples. DMN-ZS [142] introduces an attention mechanism to compute the similarity between the new test samples and the support set in order to obtain the weighted features. TDA [143] sets up two cache banks to store the one-hot labels and the negative soft labels of the reliable samples, respectively. The new test samples interact with them to generate predictions. Zhang et al. [144] employ a cache bank to save high-confidence samples to learn high-quality prompts.
In addition, MTA [145] proposes a new loss function to optimize visual features in the semantic space.
V-D Unsupervised Adaptation
Unsupervised adaptation is suitable for offline testing environments where some or all of the test data is available. Although there are no labels available, image samples can provide a priori information to help better prediction. Existing research combines different algorithms to address this problem. UPL [146] fuses prompt learning and self-supervised learning to learn prompt representations through a small number of high-quality pseudo-labels. MUST [147] incorporates image reconstruction techniques to facilitate the model’s understanding of visual information. InMaP [148] introduces prototype learning, which mines a set of visual prototypes from the visual space to replace the text embeddings for prediction. LaFTer [149] trains a text classifier and then uses it to classify images and trains the image encoder with pseudo-labels. ZLaP [150] presents a label propagation algorithm to assign labels to images. In addition, CLIPPR [151] introduces a label distribution prior to predict the data of target domain.
VI Technical Details
TABLE I: A list of commonly used benchmark datasets for zero-shot image recognition.
Dataset Year Images Class/Attr. Download
Zero-Shot Object Recognition
CUB [9] 2011 11,788 200/312 [Link]
Flowers102 [152] 2008 8,189 102/- [Link]
SUN [10] 2012 14,340 717/102 [Link]
NABirds [153] 2015 48,562 1011/- [Link]
DeepFashion [41] 2016 289,222 46/1000 [Link]
AWA [154] 2013 30,475 50/85 [Link]
AWA2 [4] 2018 37,322 50/85 [Link]
APY [155] 2009 15,339 32/64 [Link]
Zero-Shot Compositional Recognition
UT-Zappos [156] 2014 50,025 12/16 [Link]
MIT-States [157] 2015 53,753 245/115 [Link]
C-GQA [111] 2021 39,298 674/413 [Link]
AO-CLEVr [42] 2020 180,000 3/8 [Link]
VAW-CZSL [158] 2022 92,583 541/440 [Link]
Clothing16K 2020 16,170 8/9 [Link]
Foundation model-Based Open-world Recognition
ImageNet [1] 2009 1,431,167 1000 [Link]
Caltech101 [159] 2004 9,145 101 [Link]
Caltech256 [160] 2007 20,607 256 [Link]
SUN397 [161] 2010 108,753 397 [Link]
Food101 [162] 2014 101,000 101 [Link]
Flowers102 [152] 2008 8,189 102 [Link]
StanfordCars [163] 2013 16,185 197 [Link]
FGVCAircraft [164] 2013 10,000 100 [Link]
OxfordPets [165] 2012 7,349 37 [Link]
DTD [166] 2014 5,640 47 [Link]
EuroSAT [167] 2019 27,000 10 [Link]
Country211 [29] 2021 64,300 211 [Link]
CIFAR-10 [168] 2009 60,000 10 [Link]
CIFAR-100 [168] 2009 60,000 100 [Link]
Birdsnap [169] 2014 49,829 500 [Link]
[Class/Attr.] means the number of classes and attributes for ZSOR, objects and states for ZSCR, classes for FBOR.
TABLE II: More details on selected advances for zero-shot object recognition and compositional recognition.
Topic Tax. Name Venue Year Backbone Aux. KeyWords Dataset Code
Zero-Shot Object Recognition Visual LH2B [59] CVPR 2017 VGG16 Region Detection CUB, NAB [Link]
S2GA [58] NeurIPS 2018 VGG16 Region Detection CUB, NAB [Link]
AGAA [60] CVPR 2018 VGG16 Region Detection CUB, NAB [Link]
LDF [61] CVPR 2018 GNet, VGG19 - Crop CUB, AWA [Link]
AREN [36] CVPR 2019 ResNet101 - Attention CUB, SUN,AWA2, APY [Link]
SGMA [35] NeurIPS 2019 VGG19 - Crop CUB, FLO, AWA [Link]
RGEN [37] ECCV 2020 ResNet101 - Attention, Graph CUB, SUN, AwA2, APY -
RSAN [57] CIKM 2021 ResNet101 - Attention CUB, SUN, AWA2 -
SR2E [62] AAAI 2021 ResNet101 - Crop CUB, SUN, AWA2, APY -
VGSE-SMO [67] CVPR 2022 ResNet50 - Crop, Clustering CUB, SUN, AWA2 [Link]
ERPCNet [63] TCSVT 2022 ResNet101 - Crop CUB, SUN, AWA2, APY -
GKU [38] AAAI 2023 ResNet34 Region Graph CUB, NAB -
EOPA [170] TPAMI 2023 ANet, ResNet50 - Graph CUB, SUN, FLO, AWA2 -
POPRNet [64] TIP 2024 ResNet101 - Crop CUB, SUN, AWA2 [Link]
ZSLViT [65] CVPR 2024 ViT-B - Attention CUB, SUN, AWA2 -
Visual-Attribute DAZLE [39] CVPR 2020 ResNet101 - Attention CUB, SUN, AWA2, DEP [Link]
Composer [74] NeurIPS 2020 ResNet101 - Attention, Generation CUB, SUN, AWA2, DEP [Link]
APN [83] NeurIPS 2020 ResNet101 - Prototype CUB, SUN, AWA2 [Link]
GEM [73] CVPR 2021 ResNet101 Region Attention CUB, SUN, AWA2 [Link]
DPPN [85] NeurIPS 2021 ResNet101 - Prototype CUB, SUN, AWA2, APY [Link]
LAPE [171] TCSVT 2022 ResNet101 - Attention CUB, SUN, AWA2, APY -
RIAE [70] IS 2022 ResNet101 - Attention, Graph CUB, SUN, AWA2 -
DPDN [48] MM 2022 ResNet101 - Attention, Prototype CUB, SUN, AWA2 -
GNDAN [71] TNNLS 2022 ResNet101 - Attention, Graph CUB, SUN, AWA2 [Link]
AREES [75] TNNLS 2022 ResNet101 - Attention, Generation CUB, SUN, AWA, AWA2, APY -
MSDN [79] CVPR 2022 ResNet101 - Attention CUB, SUN, AWA2 [Link]
TransZero [40] AAAI 2022 ResNet101 - Attention CUB, SUN, AWA2 [Link]
I2DFormer [76] NeurIPS 2022 ViT-B - Attention CUB, FLO, AWA2 [Link]
DUET [77] AAAI 2023 ViT-B - Attention, Contrast CUB, SUN, AWA2 [Link]
PSVMA [78] CVPR 2023 ViT-B - Attention CUB, SUN, AWA2 [Link]
HRT [172] PR 2023 ResNet101 - Attention CUB, SUN, AWA2 -
CC-ZSL [87] TCSVT 2023 ResNet101 - Prototype, Contrast CUB, SUN, AWA2 [Link]
CoAR-ZSL [88] TNNLS 2023 R101, ViT-L - Attention, Prototype CUB, SUN, AWA2 [Link]
GIRL [86] IS 2023 ResNet101 - Prototype CUB, SUN, AWA2, APY [Link]
Attribute MCZSL [91] CVPR 2016 VGG16 Region Crop CUB -
LFGAA [89] ICCV 2019 GNet, R101, V19 - Attention CUB, SUN, AWA2 [Link]
APNet [90] AAAI 2020 ResNet101 - Graph CUB, SUN, AWA, AWA2, APY -
DEDN [92] IJCAI 2024 ResNet101 - Attention, Distillation CUB, SUN, AWA2 [Link]
Zero-Shot Compositional Recognition Decoupling Nan et al. [93] AAAI 2019 ResNet18 - Contrast UTZ, MIT -
CAUSAL [42] NeurIPS 2020 ResNet18 - Mapping Function UTZ, AOC [Link]
ProtoProp [94] NeurIPS 2021 ResNet18 - Prototype, Graph UTZ, AOC, CGQ [Link]
OADis [98] CVPR 2022 ResNet18 - Attention UTZ, MIT, VAW [Link]
Zhang et al. [101] ECCV 2022 ResNet18 - Attention UTZ, AOC, CLO [Link]
DECA [97] TMM 2022 ResNet18 - Mapping UTZ, MIT [Link]
DRANet [173] ICCV 2023 ResNet18 - Attention UTZ, MIT, CGQ -
ADE [99] CVPR 2023 ViT-B - Attention UTZ, CGQ, CLO [Link]
DFSP [102] CVPR 2023 CLIP - Attention, Prototype UTZ, MIT, CGQ [Link]
Hu et al. [95] AAAI 2023 ResNet18 - Contrast UTZ, CGQ [Link]
CAILA [96] WACV 2024 CLIP - Mapping UTZ, MIT, CGQ, VAW [Link]
Jing et al. [100] AAAI 2024 CLIP - Attention UTZ, MIT, CGQ -
Troika [174] CVPR 2024 CLIP - Attention UTZ, MIT, CGQ [Link]
Dependency Nagarajan et al. [8] ECCV 2018 ResNet18 - Contrast UTZ, MIT [Link]
SymNet [105] CVPR 2020 ResNet18 - Mapping Function UTZ, MIT [Link]
SAD-SP [107] TPAMI 2023 ResNet18 - Mapping Function UTZ, MIT, CGQ -
CoT [103] ICCV 2023 ResNet18, ViT-B - Attention MIT, CGQ, VAW [Link]
CANet [106] CVPR 2023 ResNet18 - Mapping Function UTZ, MIT, CGQ [Link]
Panda et al. [104] PR 2024 ResNet18 - Graph UTZ, MIT, CGQ -
ProCC [108] AAAI 2024 ResNet18 - Mapping Function UTZ, MIT, CGQ [Link]
Relationship Wei et al. [49] ICCV 2019 ResNet18 - Contrast UTZ, MIT -
BMPNet [109] TMM 2021 ResNet18 - Attention UTZ, MIT [Link]
CGE [111] CVPR 2021 ResNet18 - Graph UTZ, MIT, CGQ [Link]
SCEN [115] CVPR 2022 ResNet18 - Contrast UTZ, MIT, CGQ [Link]
CVGAE [112] MM 2022 ResNet18 - Graph UTZ, MIT, CGQ [Link]
CAPE [110] WACV 2023 ResNet18 - Attention UTZ, MIT, CGQ -
Yang et al. [116] TMM 2023 ResNet18 - contrast UTZ, MIT
GIPCOL [113] WACV 2024 CLIP - Graph UTZ, MIT, CGQ [Link]
Others TMN [5] ICCV 2019 ResNet18 - Gated Network UTZ, MIT [Link]
MUST [118] PR 2024 ResNet18 - Weighted Loss UTZ, MIT, CGQ [Link]
ProLT [119] AAAI 2024 ResNet18, CLIP - Weighted Loss UTZ, MIT, CGQ -
CDS-CZSL [120] CVPR 2024 CLIP - Weighted Loss UTZ, MIT, CGQ -
[Tax.] means taxonomy. [Aux.] means auxiliary. [GNet] means GoogleNet, [ANet] means AlexNet, [R101] means ResNet101, [V19] means VGG19.
Zero-Shot Object Recognition: [Region] means region-level annotations. [FLO] means Flowers102, [NAB] means NABirds, [DEP] means DeepFashion.
Zero-Shot Compositional Recognition: [UTZ] means UT-Zsppos, [MIT] means MIT-States, [CGQ] means C-GQA, [AOC] means AO-CLEVr, [VAW] means VAW-CZSL, [CLO] means Clothing16K.
TABLE III: More details on selected advances for fundation model-based open-world recognition.
Topic Name Venue Year VEncoder TEncoder Auxiliary Contribution Code
ZSP Menon et al. [33] arXiv 2022 ViT-B, ViT-L Transformer GPT-3 Generate description words according to category names. -
CuPL [134] ICCV 2023 ViT-L/14 Transformer GPT-3 Generate description sentences according to category names. [Link]
SuS-X [136] ICCV 2023 ResNet50 Transformer Diffusion Generate visual support set with Diffusion. [Link]
CHiLS [135] ICML 2023 ViTL/14@336px Transformer GPT-3 Generate sub-class names according to category names. [Link]
CALIP [34] AAAI 2023 ResNet50 Transformer - Propose attention interaction of image and text embeddings. [Link]
TTA TPT [132] NeurIPS 2022 ResNet50, ViT-B/16 Transformer - Propose test-time prompt adaptation. [Link]
DiffTPT [138] ICCV 2023 ResNet50, ViT-B/16 Transformer Diffusion Propose to augment image with Diffusion. [Link]
PromptAlign [139] NeurIPS 2023 ViT-B/16, ViT-B/32 Transformer Statistic Align the distributions of test and pre-training set. [Link]
SwapPrompt [140] NeurIPS 2023 ResNet50 Transformer - Update the prompt using exponential moving average. -
Zhang et al. [144] AAAI 2024 ViT-B/16 Transformer - Introduce multiple manual prompts. -
DART [141] AAAI 2024 ViT-B/16 Transformer - Learn a separate prompt for each class. -
TDA [143] CVPR 2024 ResNet50, ViT-B/16 Transformer - Two cache banks are set up to store hard and soft labels. [Link]
MTA [145] CVPR 2024 ViT-B/16 Transformer - Improve visual representation in the embedding space. [Link]
DMN-ZS [142] CVPR 2024 ResNet50, ViT-B/16 Transformer - Cache high-confidence samples for subsequent predictions. [Link]
UA CLIPPR [151] arXiv 2022 ViT-B/32 Transformer Distribution Predicted label aligns a priori label distribution. [Link]
UPL [146] arXiv 2022 ResNet50 Transformer - Select high-confidence samples for fine-tuning. [Link]
MUST [147] ICLR 2023 ViT-B/16, ViT-L/14 Transformer - Introduce visual reconstruction and teacher-student model. [Link]
InMaP [148] NeurIPS 2023 ResNet50, ViT-B/16 Transformer - Learn visual prototypes in the embedding space. -
LaFTer [149] NeurIPS 2023 ViT-B/32 Transformer GPT-3 Train a text classifier to categorize images. -
ZLaP [150] CVPR 2024 ResNet50, ViT-B/16 Transformer - Introduce a label propagation algorithm. [Link]
[ZSP] means zero-shot prediction. [TTA] means test-time adaptation. [UA] means unsupervised adaptation. [VEncoder] means visual encoder. [TEncoder] means text encoder. [ViT-B]: ViT-B/32, ViT-B/16. [ViT-L]: ViT-L/14, ViT-L/14@336px. [Statistic] means the mean and variance of pretrained dataset. [Distribution] means the label distribution.
VI-A Dataset
We list in Table I the image recognition evaluation datasets commonly used for the three tasks. There are 8 datasets used for zero-shot object recognition, including fine-grained datasets such as CUB [9] which is a bird dataset, SUN [10] which is a scene dataset, and coarse-grained datasets such as AWA [154] which is a dataset of various types of animals. Most of the datasets contain detailed attribute annotations. Notice that Flowers102 [152] and NABirds [153] do not have attribute annotations. Flowers102 contains textual descriptions, i.e., each image is described by several sentences. NABirds has region annotations, which mark the parts of the bird, and are typically used in visual component analysis methods.
There are 6 commonly used datasets for zero-shot compositional recognition. The columen [Class/Attr.] indicates the number of objects and the number of states they contain. It can be seen that their compositional sizes cover different orders of magnitude, ranging from $3/8$ (AO-CLEVr [42]), which means up to 24 compositions, to $674/413$ (C-GQA [111]), which means 278,362 compositions.
We show 15 commonly used evaluation datasets for foundation model-based open-world recognition that cover different categories and different data sizes. ImageNet [1] includes data with up to 1000 categories, while EuroSAT [167] and CIFAR-10 [168] have only 10 categories. The data types are also diverse, including food [162], flowers [152], pets [165], transportation [163, 164], etc. They are also fine-grained datasets. In addition, DTD [166] is a texture dataset and EuroSAT [167] is a remote sensing dataset.
VI-B Implement Details
To facilitate understanding and referring readers to related work, we provide more details of the latest developments in Tables II and III. Table II shows the implementation details of the related studies of ZSOR and ZSCR, including the backbone, auxiliary, dataset, and code addresses. It can be seen that commonly used backbones include ResNet [175], VGG [176], GoogleNet [177], AlexNet [178], ViT [66], and CLIP [29]. [Auxiliary] refers to additional information such as region-level annotations. We provide the dataset and code address to facilitate access to the performance of the method as well as its implementation. Finally, we label the study with up to two keywords based on the specific technique used.
Table III shows more details of the related work in foundation model-based open-world recognition. We list their visual and text encoders, as well as auxiliary information. GPT-3 [133] and Diffusion [137] are the exotic models. [Statistic] denotes the mean and variance of the pre-training data. [Distribution] denotes the prior distribution of labels. We also provide their code addresses for quick reference. Since FBOR is at a young stage with less related work, it is not easy to categorize. To better understand the focus of each work, we summarize their contributions in a single sentence.
VI-C Shared Technologies
Observing Tables II and III, we can identify some techniques that are common across the three tasks. We hope to discuss these techniques to reveal more commonalities between several tasks. Attention is the most commonly used technique, distributed across almost every sub-direction in the three tasks. This suggests that attention techniques are extremely generalizable and can be applied in different task goals and with different research motivations. In ZSOR, attention is used to localize key visual regions [36, 37], model visual-text correlations [39, 79, 40], and localize key attributes [89]. In ZSCR, attention is used to model object and state correlations [98, 101, 103], and enhance representations [109, 110]. In FBOR, attention is used to capture correlations between images and text [34]. Prototype learning is another more generalized technique. In ZSOR, prototypes are used to represent attributes [83, 85]. In ZSCR, prototypes are used to represent objects and states [94, 102]. In FBOR, prototypes are used to replace the embedding of prompts in deep space [148]. Graph and contrast learning are the more commonly used techniques in ZSOR and ZSCR. Graph is used to model the topological relationships between visual regions in ZSOR [70, 38], while in ZSCR it is used to capture the logical relationships between individual objects and states [111, 112]. Contrast learning is used in ZSOR to optimize attribute embeddings [77], while in ZSCR it is used to untangle and model semantic relations [49, 115]. In FBOR, thanks to the robust zero-shot performance of VLMs, fine-tuning prompts becomes a popular approach [132, 140]. Similarly, in ZSCR, some methods with CLIP as the backbone introduce prompt learning [102, 113], which effectively improves the performance of the model.
VII Application
Zero-shot learning is of great practical significance due to its radical freeing up of the pressure of image data collection. Meanwhile, element-wise zero-shot learning provides an intuitive humanoid study pattern which is suitable for a variety of reasoning tasks. In recent years, besides image recognition, element-wise zero-shot techniques have been applied in a wide range of areas, and they promote each other and develop synergistically.
Language and Text. Language, as one of the main media for human beings, has always received as much attention from researchers as images. Language research is also under pressure to label data. Therefore, zero-shot generalization ability is a long standing direction of attack [179, 180, 181]. In recent years, large language models have been developed rapidly [182], and they can even talk with humans fluently. This shows that large language models have powerful zero-shot reasoning ability. Meanwhile, some studies [183] propose fine-tuning schemes for language models to improve the zero-shot performance. In some specific tasks, zero-shot techniques have been widely applied to solve the problem of insufficient data samples, such as machine translation [184], question answering [185], etc. In addition, zero-shot techniques also play an important role in text analysis. Text analysis is the classification of text, such as sentiment classification [186], topic classification [187, 188], etc., which is extremely helpful for research in sociology, psychology and other disciplines. The content of text is quite complex, and the text elaborating the same topic may have multiple forms of expression. Obviously, the collection and labeling of samples is a huge obstacle, and zero-shot learning offers a promising solution for it. Moreover, zero-shot learning has been intensively studied in various fields such as named entity recognition [189] and information retrieval [190].
Scene Understanding. Scene understanding is a topic based on image recognition but more challenging. Instead of assigning a label to an entire image, it usually identifies and localizes multiple elements in an image. For example, multi-label recognition task [191, 192] requires simultaneous recognition of multiple objects in an image. In contrast, object detection [193, 194, 195] and semantic segmentation [196, 197, 198] are two more complex and practical tasks. Object detection requires identifying and localizing multiple objects in an image, while semantic segmentation is a pixel-level classification task that requires assigning labels to each pixel according to semantic information. Scene understanding is an important research topic, which is not only the basis for AI to understand the world, but also a key part of deployment to industrial scenarios. Related technologies are needed in a variety of scenarios such as autonomous driving, medical diagnostic assistance, security monitoring, retail, and so on. However, they also require more detailed annotation and larger data for training than image recognition. Therefore, incorporating zero-shot learning into scene understanding has become an attractive direction in recent years. For example, zero-shot object detection [199, 200, 201] and zero-shot semantic segmentation [202, 203, 204] are dedicated to the deployment of models in the open world.
3D Vision. With the increasing maturity of 3D sensors, the cost of acquiring 3D vision data has dropped significantly [205]. Compared with traditional 2D images, 3D visual data contains important depth information, and therefore can provide richer semantic information. 3D vision data has a wide range of applications, covering automatic driving, medical diagnosis, remote sensing, etc. Correspondingly, research on 3D vision data has also received more and more attention. However, compared with 2D data, the acquisition and labeling cost of 3D data is more expensive. Therefore, many researches integrate the concept of zero-shot into 3D tasks, such as point cloud classification [206, 207, 208], object detection [209], and semantic segmentation [210, 211]. In addition to semantic recognition tasks, some studies are also exploring the application of zero-shot learning to tasks such as generation [212], style transfer [213], and synthesis [214, 215]. There are also some studies that have explored in-depth on special tasks, such as human pose estimation [216], shape correspondence [217], and so on.
Other Tasks. Zero-shot learning offers promising solutions for many specific application scenarios. For example, medical and remote sensing images are much more difficult to collect than natural images due to sample scarcity and data protection. Therefore, the research on zero-shot medical image [218, 219] recognition and remote sensing image [220, 221] recognition has been developed rapidly in recent years. Anomaly detection [222] plays a key role in securing many industrial fields. However, the number and type of anomaly samples are often insufficient. Some studies resort to zero-shot learning to solve this problem. In addition, zero-shot techniques have been applied in many special tasks to alleviate the low-data problem as well as to enhance the generalization ability of the model, such as action recognition [223], image retrieval [224, 225], and video classification [226].
VIII Future Directions
Zero-shot learning contributes powerful technical solutions for automated recognition in the open environment. Meanwhile, research based on element-wise representation and reasoning has driven the rapid development and wide application of zero-shot technology. Despite the remarkable achievements, there are still some potential limitations that have not been deeply explored. In this section, we share our perspectives on future research directions.
VIII-1 Continual Zero-Shot Learning
Existing studies are based on closed systems to design solutions. However, in reality there are a large number of growing systems with dynamically changing environments and goals. For example, for zero-shot object recognition, attributes and categories may be progressively inflated. Early in the observation phase, fewer attributes are accessible, followed by a steady stream of new attribute annotations added. Similarly, the detected categories may increase over time forward. It is difficult for existing schemes to handle such dynamic recognition tasks, and continual learning has come a long way in this regard [227]. However, simply transplanting techniques from continual learning to zero-shot recognition tasks suffers from deviation. Therefore, fusing continual learning and zero-shot learning will further increase the value of applications to real-world scenarios.
VIII-2 Multi-label Compositional Recognition
Zero-shot compositional recognition allows the model to not only learn to recognize objects, but also to describe their state. However, current research is limited to single object and single state. A robust compositional recognition system should have the ability to recognize multiple objects as well as multiple states. For a single object, the recognition system should describe its state from multiple perspectives to provide more detailed information. For multiple objects, their individual states as well as their interactions, relationships, etc. should be described. Expanding from single label to multiple labels is a tough choice, but it is an inescapable step towards building a smarter recognition system.
VIII-3 Train VLM with Downstream Fine-Grained Annotations
The power of VLMs in the open world has been demonstrated by the fact that it requires only the names of the categories in the downstream task to accomplish zero-shot inference. However, the performance of VLMs in tasks with fine-grained annotations has not been explored. If the downstream tasks have more refined element-wise annotations, such as attribute texts or state descriptions, can VLMs fully utilize this information to obtain better zero-shot recognition performance. There are a number of issues that deserve deeper investigation, including domain shift between pre-trained text embeddings of VLMs and downstream text, fine-grained visual-text matching, catastrophic forgetting, and so on.
VIII-4 Annotation Cost and Quality
Element-wise schemes in zero-shot object recognition typically require extensive refined attribute annotations, especially for visual-attribute matching schemes. However, the attribute-level annotations are time- and labor-intensive compared to class-level labeling. Worse still, once they settle into concrete real-world scenarios, such as industrial inspection or medical pathology, the expert knowledge can be a bottleneck, which further raises the labor cost. In addition, attribute engineering is a complex crossover field. Even attributes annotated by experienced experts do not guarantee benefits for deep learning, which implies that high-quality attribute annotations require professionals with dual knowledge of both specific domains and deep learning. Despite some studies attempting to make breakthroughs in the field of automated annotation [91], it is clear that there is still a long way to go.
VIII-5 Deployment Cost
Compared to class-wise semantic modeling, many element-wise schemes have to process a higher density of information, which introduces a more luxurious deployment cost. Such cost is reflected in bloated network structures and high computational complexity (Note that we discuss the deployment phase, excluding the training phase). As a result, they are unfriendly to edge tasks and mini-endpoints, which have to trade off performance and memory. However, they can be more favorable to a scenario associated with resource-constrained devices due to the low or even zero data requirements. Such a scenario can also well align with ubiquitous devices and data in real-world applications. Therefore, it is promising to investigate on-device-friendly element-wise algorithms.
VIII-6 Unified Framework
Currently, zero-shot research for all directions is booming. They share many techniques such as attention, graph modeling, and so on. However, barriers between the them still exist due to the different goals and contexts of the tasks. A technique and framework that is versatile across all fields is lacking. But the significance of such a unified framework for both academia and application is immense. The quest for a unified framework means exploring more commonalities across zero-shot tasks, and thinking about the operational mechanisms of the entire realm from the ground-up principles. From a realistic point of view, a mature intelligence should be able to handle multiple tasks at the same time.
VIII-7 Robust Representation Learning
Domain shift is a crucial hindrance to the generalization of zero-shot techniques to the open world, and one of its solutions is to learn robust feature representation. Although many methods have been proposed to mitigate the effects of domain shift, quantitative tests of the resistance to interference are lacking, and thus the robustness of existing representation methods in real-world environments remains unknowable. Recently Shafiee et al. [228] introduce adversarial attacks into zero-shot models and propose a framework for combining attributes to generate perturbations that effectively degrade the performance of zero-shot models. Attack and defense of models are two topics that have coexisted for a long time. Attackers seek to weaken the model at minimal cost, while defenders need to resist attacks. They learn from each other and progress continuously. Very often, the defender must learn more robust representation in order to resist perturbations. Thus, for zero-shot learning, it provides a new perspective to combat domain shift.
VIII-8 Poor Theoretical Foundation
The development of zero-shot learning is established on the beautiful hypothesis that deep neural networks can reason logically like humans. As mentioned earlier, we can learn to identify basic elements and recognize new concepts by combining and reasoning about them. Nevertheless, there are not many solid theories on the compatibility between human reasoning and machine inductive ability so far, leading to a lack of explainability. Meanwhile, some flaws also challenge the plausibility of the hypothesis, such as the correspondence between abstract attributes and vision. Rigorous theoretical guidance is at the helm of a field moving forward, and it is of great prospective to dive into the mysterious black box in the future.
IX Conclusion
We provide a comprehensive exposition of three major tasks in the field of zero-shot image recognition that share the notion of element-wise representation and reasoning. Specifically, we show the background, commonalities, and challenges faced by several tasks. Then related recent techniques are meticulously categorized and explained. At the same time, we provide specific technical details, including datasets, training details, code repository addresses, and more. Finally, we show the wide range of applications of element-wise zero-shot learning and look at future directions.